How much water must be evaporated form 340 pounds of a solution that is 3% alcohol to strengthen it to a solution that is 5% alcohol
Answer by stanbon(75887)

How much water must be evaporated form 340 pounds of a solution that is 3% alcohol to strengthen it to a solution that is 5% alcohol
---------
Equation::
water - water = water
0.97*340 - 1*x = 0.95(340-x)
-------
97*340 - 100x = 95*340 - 95x
5x = 2*340
x = 2*68 = 136 lbs (amt. of water to evaporate)
-----------------
Cheers,
Stan H.
================